OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of −3 and passes through the point (−2,1

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

y = -3x + 1, the formula is y = mx + b

OpenStudy (anonymous):

y – y1 = m(x – x1) y – (–6) = (4)(x – (–1)) y + 6 = 4(x + 1) y + 6 = 4x + 4 y = 4x + 4 – 6 y = 4x – 2

OpenStudy (anonymous):

b is the y intercept, in this case, 1, and mx is the slope, in this case, -3

OpenStudy (anonymous):

If you don't mind, let me try to help. Let y=mx + p, the equation of the line. m is the slope. In this case, m=-3 . p is the y-intercept. It's unknown. But, since the line passes trough the point(-2, 1), we can find the value of P by substitute the values of the point into the equation. where x=-2 and y=1. we have: y = -3x + p with the substitution, it becomes: 1=-3(-2)+p or 1=6+p or p=1-6 or p=-5 Therefore, the equation of the line is: y = -3x - 5 I wish I have helped you.

OpenStudy (anonymous):

Even with the expression used by ameenbriscoe13, the result will be y= -3x-5, not y = 4x-2. But you just have to substitute y1 by 1 , m by -3 and x1 by -2.